# Sequential Regression Learning with Randomized Algorithms

### v1

@@ 3. Absolutely Continuous Random Algorithm
The Equation 3.3 tells us that
$c_{(x_{t},y_{t})}(g_{t})+\frac{1}{2\eta_{t-1}}\parallel g_{t}-g_{t-1}\parallel_{L^{2}}^{2}=\Theta_{(x_{t},y_{t})}(g_{t})+\frac{1}{2\eta_{t-1}}\parallel g_{t}-g_{t-1}\parallel_{L^{2}}^{2}\leq\Theta_{(x_{t},y_{t})}(g_{t-1})=c_{(x_{t},y_{t})}(g_{t-1}),$
for each $t\in\mathbb{N}$. As $\Theta_{(x_{t},y_{t})}$ is convex, it follows from Equation 3.4 that
$\Theta_{(x_{t},y_{t})}(g_{t})-\Theta_{(x_{t},y_{t})}(g^{\star})\leq\langle\nabla\Theta_{(x_{t},y_{t})}(g_{t}),g_{t}-g^{\star}\rangle_{L^{2}}=\langle-\frac{g_{t}-g_{t-1}}{\eta_{t-1}},g_{t}-g^{\star}\rangle_{L^{2}}.$
Moreover, it follows from parallelogram law that
$2\langle-(g_{t}-g_{t-1}),g_{t}-g^{\star}\rangle_{L^{2}}$ $=$ $2\langle g_{t-1}-g_{t},g_{t}-g^{\star}\rangle_{L^{2}}=\parallel g_{t-1}-g_{t}+g_{t}-g^{\star}\parallel^{2}-\parallel g_{t-1}-g_{t}\parallel^{2}$
-- $\parallel g_{t}-g^{\star}\parallel^{2}~{}\leq~{}\parallel g_{t-1}-g^{\star}\parallel^{2}-\parallel g_{t}-g^{\star}\parallel^{2}.$
As a consequence
$2\eta_{t-1}\left[c_{(x_{t},y_{t})}(g_{t})-c_{(x_{t},y_{t})}(g^{\star})\right]$ $=$ $2\eta_{t-1}\left[\Theta_{(x_{t},y_{t})}(g_{t})-\Theta_{(x_{t},y_{t})}(g^{\star})\right]$
$\leq$ $~{}2\langle g_{t-1}-g_{t},g_{t}-g^{\star}\rangle_{L^{2}}~{}\leq~{}\parallel g_{t-1}-g^{\star}\parallel_{L^{2}}-\parallel g_{t}-g^{\star}\parallel_{L^{2}}.$
Summing up, we obtain
$2\sum_{t=2}^{n+1}\eta_{t-1}\left[c_{(x_{t},y_{t})}(g_{t})-c_{(x_{t},y_{t})}(g^{\star})\right]\leq\parallel g_{1}-g^{\star}\parallel_{L^{2}}-\parallel g_{n+1}-g^{\star}\parallel_{L^{2}}\leq\parallel g_{1}-g^{\star}\parallel_{L^{2}}$
for every $n\geq 1$. ∎
In the sequel, we will use Lemma 3.1 to establish an online learning algorithm. We set $A_{1}(dh)=g_{1}(h)\delta(dh)$ with $g_{1}=1$. Then, we observe $(x_{1},y_{1})\in\mathbb{O}$ and update the learner’s action as follows:
$g_{2}(\cdot\mid{\bf o}_{1})=\operatorname*{prox}_{\eta_{1}\Psi}\left[g_{1}-\eta_{1}\nabla c_{(x_{1},y_{1})}(g_{1})\right]\quad\text{and}\quad A_{2}(dh\mid{\bf o}_{1})=g_{2}(h\mid{\bf o}_{1})\delta(dh).$
For any $t\geq 3$, we observe $(x_{t},y_{t})\in\mathbb{O}$ and update the learner’s action as follows:
(3.5) $g_{t+1}(\cdot\mid{\bf o}_{t})=\operatorname*{prox}_{\eta_{t}\Psi}\left[g_{t}(\cdot\mid{\bf o}_{t-1})-\eta_{t}\nabla c_{(x_{t},y_{t})}(g_{t})\right]~{}~{}\text{and}~{}~{}A_{t+1}(dh\mid{\bf o}_{t})=g_{t+1}(h\mid{\bf o}_{t})\delta(dh).$
Theorem 3.1.
The sequence of proximal algorithm $\{A_{t}:t\geq 1\}$ given by 3.5 is PAC. In this case, we have
$\frac{1}{n}\sum_{t=2}^{n}\eta_{t-1}\left[\int_{\mathcal{H}}\ell(y_{t},h(x_{t}))A_{t}(dh\mid{\bf o}_{t-1})\delta(dh)-\int_{\mathcal{H}}\ell(y_{t},h(x_{t}))A^{\star}(dh)\delta(dh)\right]\leq\frac{1}{2n}\parallel g_{1}-g^{\star}\parallel_{L^{2}},$
where $A^{\star}(dh)=g^{\star}(h)\delta(dh)$ and $n\geq 2$.
If we take ${\bf o}_{t-1}=(x_{1},y_{1}),\cdots,(x_{t-1},y_{t-1})\in\mathbb{O}^{t-1}$ and $x\in\chi$, given $\mathcal{H}$ a Borel space and $\delta$ a probability on $(\mathcal{H},\beta(\mathcal{H}))$, the estimate $\hat{y}_{x}$ is a probability distribution over $(\mathcal{Y},\beta(\mathcal{Y}))$ defined by (see, Equation 3.5)
(3.6) $\hat{y}_{x}(t,F)$ ::=:= $\int_{\mathcal{H}}1\!\!1_{\{F\}}(h(x))A_{t}(dh\mid{\bf o}_{t-1})=\int_{\mathcal{H}}1\!\!1_{\{F\}}(h(x))g_{t}(h)\delta(dh),$
where $F\in\beta(\mathcal{Y})$. If we take $\mathcal{Y}:=\mathbb{R}^{p}$ with $p\geq 1$, we can predict the output of the process using the following conditional expectation:
$\hat{y}_{x}(t):=\int_{\mathbb{R}^{p}}y~{}\hat{y}_{x}(t,dy),\quad t=2,3,\cdots.$
4. Randomized SINDy
Consider data measurements $y\in\mathbb{R}^{m}$ that may be approximated by a linear combination of columns from a feature library $\Delta(x)\in\mathbb{R}^{m\times p}$, for every $x\in\mathbb{R}^{k}$. The linear combination of columns is given by the entries in the random vector $\beta\in\mathbb{R}^{p}$ so that
$Y=\Delta(x)\beta+\epsilon$
where $\epsilon$ represents the random error. In this case, we take
$\Delta(x)=\left(\Delta_{1}(x),\Delta_{2}(x),\cdots,\Delta_{d}(x)\right)$
We suppose that $\beta$ has a $d$-dimensional multivariate distribution with density $g_{b}$ parametrized by $b\in\mathcal{B}\subset\mathbb{R}^{w}$. Thus, we can identify strategies as the parameters $(\mu,\Sigma)$ of the multivariate normal distribution. Then, for every pair of $o=(x,y)\in\mathbb{O}$, we take
(4.1) $c_{(x,y)}(b):=\int_{\mathbb{R}^{p}}\ell(y,\Delta(x)\beta)g_{b}(\beta))d\beta=\mathbb{E}_{b}\ell(y,\Delta(x)\beta)$
where $b\in\mathcal{B}$. We consider the following stochastic Ridge regression:
(4.2) $\hat{b}=arg\min_{b}\left\{\mathbb{E}_{b}\ell(y,\Delta(x)\beta)+\lambda\mathbb{E}_{b}\beta^{T}\beta+\Psi(b)\right\}$
where $\lambda\geq 0$ and
$\Psi(b)=\begin{cases}0&;~{}~{}~{}b\in b\in\mathcal{B}\\ \infty&;~{}~{}~{}g\not\in b\in\mathcal{B}.\end{cases}$
Example 1
Considering a multiple linear regression model $y=\mathbf{x}^{\top}\beta+\epsilon$ where $\mathbf{x}=\Delta(x)$ and $\beta=\left(\beta_{1},\cdots,\beta_{p}\right)$ are both vectors of dimension $p\times 1$ and $\epsilon\sim\mathcal{N}(0,\sigma^{2})$, the space of predictors $\mathcal{H}$ consists now of all possible linear predictors on the $p$-dimensional space. A linear predictor is defined by a weight vector $\mathbf{w}\in\mathbb{R}^{p}$ such that $h(\mathbf{x})=\mathbf{x}^{\top}\mathbf{w}$.
@@ 4. Randomized SINDy
We assume the probability density over the weight vector $\mathbf{w}$ follows a multivariate normal distribution $\mathcal{N}(\mu,\Sigma)$ with mean vector $\mu=(\mu_{1},\cdots,\mu_{p})$ and covariance matrix $\Sigma=\operatorname{diag}(\sigma_{1},\cdots,\sigma_{p})$:
$g(\mathbf{w})=\frac{1}{(2\pi)^{p/2}|\Sigma|^{1/2}}\exp\left\{-\frac{1}{2}(\mathbf{w}-\mu)^{\top}\Sigma^{-1}(\mathbf{w}-\mu)\right\}.$
@@
The hypothesis space can then be re-parametrized by $\theta=\left(\mu,\Sigma\right)$, where $\mu\in\mathbb{R}^{p}$ and $\Sigma\in\mathbb{R}^{p\times p}$ (p.s.d.). By assuming this parametric form for $g$, the constraints for a probability density are inherently satisfied. At every time step, the density function $g(\mathbf{w})$ is completely determined by $\theta$. The expected loss according to Equation 4.1, is given by:
[…]
$=$ $\mathbb{E}_{W\sim\mathcal{N}(\mu,\Sigma)}~{}\ell(y,\mathbf{x}^{\top}W).$
By using the quadratic loss the optimization problem, according to Equation 4.2, is as follows:
@@
(4.3) $\operatorname*{arg\,min}_{\mu,\Sigma}\left\{~{}\mathbb{E}~{}(y-\mathbf{x}^{\top}W)^{2}+\lambda\mathbb{E}~{}W^{\top}W+\Psi(\Sigma)~{}\right\}$
@@
We now compute each term in the objective function 4.3. Since $W\sim\mathcal{N}(\mu,\Sigma)$,
$(y-\mathbf{x}^{\top}W)^{2}=y^{2}-2y\mathbf{x}^{\top}W+(\mathbf{x}^{\top}W)^{2},$
@@
$\mathbb{E}(y-\mathbf{x}^{\top}W)^{2}$ $=$ $\mathbb{E}y^{2}-2y\mathbf{x}^{\top}\mathbb{E}W+\mathbb{E}(\mathbf{x}^{\top}W)^{2}.$
@@
$\mathbb{E}[W]=\mu,$
$\text{Var}(W)=\Sigma=\mathbb{E}[WW^{\top}]-\mathbb{E}[W]\mathbb{E}[W]^{\top}\implies\mathbb{E}[WW^{\top}]=\Sigma+\mu\mu^{\top}.$
@@
$\mathbb{E}~{}(\mathbf{x}^{\top}W)^{2}$ $=$ $\mathbb{E}~{}\mathbf{x}^{\top}WW^{\top}\mathbf{x}$
$=$ $\mathbf{x}^{\top}\mathbb{E}[WW^{\top}]\mathbf{x}$
$=$ $\mathbf{x}^{\top}(\Sigma+\mu\mu^{\top})\mathbf{x}$
$=$ $\mathbf{x}^{\top}\Sigma\mathbf{x}+\mathbf{x}^{\top}\mu\mu^{\top}\mathbf{x}$
$=$ $\mathbf{x}^{\top}\Sigma\mathbf{x}+(\mathbf{x}^{\top}\mu)^{2}.$
@@
$\mathbb{E}\left[(y-\mathbf{x}^{\top}W)^{2}\right]=y^{2}-2y\mathbf{x}^{\top}\mu+\mathbf{x}^{\top}\Sigma\mathbf{x}+(\mathbf{x}^{\top}\mu)^{2}.$
@@
$\lambda\mathbb{E}W^{\top}W=\lambda\left(\text{tr}\Sigma+\mu^{\top}\mu\right).$
And the constraint term enforces that the estimated covariance matrix of $W$ is p.s.d. which can be represented using an indicator function:
@@
$\underbrace{y^{2}-2y\mathbf{x}^{\top}\mu+\mathbf{x}^{\top}\Sigma\mathbf{x}+(\mathbf{x}^{\top}\mu)^{2}+\lambda\left(\text{tr}\Sigma+\mu^{\top}\mu\right)}_{f_{1}}+\underbrace{1\!\!1_{p.s.d}(\Sigma)}_{f_{2}}$
@@
$\nabla_{\mu}f_{1}(\mu,\Sigma)=2\mathbf{x}(\mathbf{x}^{\top}\mu-y)+2\lambda\mu,$
@@
$\nabla_{\Sigma}f_{1}(\mu,\Sigma)=\mathbf{x}\mathbf{x}^{\top}+\lambda I.$
The indicator function in $f_{2}$ is non-smooth and the proximal operator is reduced to a projection onto a p.s.d. cone, following [CVX],
@@
(4.4) $\mu_{t+1}=\mu_{t}-\eta_{t}\nabla_{\mu_{t}}f_{1}(\mu_{t},\Sigma_{t})=\mu_{t}-\eta_{t}\left(2\mathbf{x}_{t}\left(\mathbf{x}^{\top}\mu_{t}-y_{t}\right)+2\lambda\mu_{t}\right),$
$\widetilde{\Sigma}_{t+1}=\Sigma_{t}-\eta_{t}\nabla_{\Sigma_{t}}f_{1}(\mu_{t},\Sigma_{t})=\Sigma_{t}-\eta_{t}\left(\mathbf{x}_{t}\mathbf{x}^{\top}_{t}+\lambda I\right),$
(4.5) $\Sigma_{t+1}=\text{proj}_{K}\left(\widetilde{\Sigma}_{t+1}\right)$
@@
Example 2
In a binary classification setting where $Y\in\{0,1\}$ and $\mathbf{x}=\Delta(x)\in\mathbb{R}^{p}$, we can adapt the previous example by utilizing the logistic function such that $h(\mathbf{x})=\sigma(\mathbf{x}^{\top}\mathbf{w})$ represents the probability of predicting the output label $Y=1$:
$P(Y=1\mid\mathbf{x},\mathbf{w})=\sigma(\mathbf{x}^{\top}\mathbf{w})=\dfrac{1}{1+\exp\{-\mathbf{x}^{\top}\mathbf{w}\}}.$

### v2

@@ 2. Randomized SINDy
We consider the following stochastic Ridge regression:
(2.1) $\hat{b}=\arg\min_{b\in\Theta}\mathbb{E}_{b}L(b,(x,y))=\arg\min_{b}\left\{L(b,(x,y))+\Psi(b)\right\}$
where $\lambda\geq 0$, $L(b,(x,y))=\mathbb{E}_{b}\ell(y,\Delta(x)\beta)+\lambda\mathbb{E}_{b}\beta^{T}\beta$ and
$\Psi(b)=\begin{cases}0&;~~~b\in\Theta\\ \infty&;~~~b\not\in\Theta.\end{cases}$
By convexity of $L$, we have that
(2.2) $L(\hat{b}^{t}({\bf o}_{t-1}),(x_{t},y_{t}))\leq L(b,(x_{t},y_{t}))+\langle\nabla L(\hat{b}^{t}({\bf o}_{t-1}),(x,y)),\hat{b}^{t}({\bf o}_{t-1})-b\rangle,\quad b\in\Theta.$
Applying the Pythagorean theorem, we have that
$\parallel\hat{b}^{t+1}({\bf o}_{t})-b\parallel^{2}$ $=$ $\parallel prox_{\eta_{t},\Psi}\left[\hat{b}^{t}({\bf o}_{t-1})-\eta_{t}\nabla L(\hat{b}^{t},(x_{t},y_{t}))\right]-b\parallel^{2}$
$\leq$ $\parallel\hat{b}^{t}({\bf o}_{t-1})-\eta_{t}\nabla L(\hat{b}^{t},(x_{t},y_{t}))-b\parallel^{2}$
$=$ $\parallel\hat{b}^{t}({\bf o}_{t-1})-b\parallel^{2}+\eta_{t}^{2}\parallel\nabla L(\hat{b}^{t},(x_{t},y_{t}))\parallel^{2}-2\eta_{t}\langle\nabla L(\hat{b}^{t},(x_{t},y_{t})),\hat{b}^{t}({\bf o}_{t-1})-b\rangle,$
for every $b\in\Theta$. Hence, we conclude that
(2.3) $2\langle\nabla L(\hat{b}^{t},(x_{t},y_{t})),\hat{b}^{t}({\bf o}_{t-1})-b\rangle\leq\frac{\parallel\hat{b}^{t}({\bf o}_{t-1})-b\parallel^{2}-\parallel\hat{b}^{t+1}({\bf o}_{t})-b\parallel^{2}}{\eta_{t}}+\eta_{t}\parallel\nabla L(\hat{b}^{t},(x_{t},y_{t}))\parallel^{2}.$
As a consequence of equations 2.2 and 2.3, we obtain that
$L(\hat{b}^{t}({\bf o}_{t-1}),(x_{t},y_{t}))-L(b,(x_{t},y_{t}))$ $\leq$ $\frac{\parallel\hat{b}^{t}({\bf o}_{t-1})-b\parallel^{2}-\parallel\hat{b}^{t+1}({\bf o}_{t})-b\parallel^{2}}{\eta_{t}}$
$+$ $\eta_{t}\parallel\nabla L(\hat{b}^{t},(x_{t},y_{t}))\parallel^{2}.$
Summing from $1$ to $n$ with $\eta_{t}=\frac{D}{G\sqrt{t}}$ and $\frac{1}{\eta_{0}}=0$, we obtain for every $b\in\Theta$ that
$\sum_{t=1}^{n}\left(L(\hat{b}^{t}({\bf o}_{t-1}),(x_{t},y_{t}))-L(b,(x_{t},y_{t}))\right)$ $\leq$ $\sum_{t=1}^{n}\parallel\hat{b}^{t}({\bf o}_{t-1})-b\parallel^{2}\left(\frac{1}{\eta_{t}}-\frac{1}{\eta_{t-1}}\right)$
$+$ $\sum_{t=1}^{n}\eta_{t}\parallel\nabla L(\hat{b}^{t},(x_{t},y_{t}))\parallel^{2}$
$\leq$ $D^{2}\sum_{t=1}^{n}\left(\frac{1}{\eta_{t}}-\frac{1}{\eta_{t-1}}\right)+G^{2}\sum_{t=1}^{n}\eta_{t}$
$\leq$ $D^{2}\frac{1}{\eta_{n}}+DG\sum_{t=1}^{n}\frac{1}{\sqrt{t}}\leq DG\sqrt{n}+DG2\sqrt{n}\leq 3DG\sqrt{n}.$
The last inequality follows since $\sum_{t=1}^{n}\frac{1}{\sqrt{t}}\leq 2\sqrt{n}$.
If we consider the $p$-dimensional vector $\hat{\beta}_{t}$ with distribution $g_{\hat{b}_{t}}$, we can predict the output of the process as follows
$\hat{y}_{t}=\mathbb{E}_{\hat{b}_{t}}\left(\Delta(x_{t})\hat{\beta}_{t}\right),\quad t\geq 1.$
As we know the distribution of the quantity $\Delta(x_{t})\hat{\beta}_{t}$, the prediction probability interval is given by
$\mathbb{P}\left[LI_{\alpha/2}\leq\Delta(x_{t})\hat{\beta}_{t}\leq LS_{\alpha/2}\right]=1-\alpha,\quad 0<\alpha<1.$
2.1. Randomized SINDy for regression
@@
Considering a sequential regression model $y=\Delta(x)\beta$ where $\beta=\left(\beta_{1},\cdots,\beta_{p}\right)$. If we assume $\beta$ follows a multivariate normal distribution $\mathcal{N}_{p}(\mu,\Sigma)$ with mean vector $\mu=(\mu_{1},\cdots,\mu_{p})$, covariance matrix $\Sigma=\operatorname{diag}(\sigma_{1},\cdots,\sigma_{p})$ and density function given by:
$g_{b}(\beta)=\frac{1}{(2\pi)^{p/2}|\Sigma|^{1/2}}\exp\left\{-\frac{1}{2}(\beta-\mu)^{\top}\Sigma^{-1}(\beta-\mu)\right\}.$
@@
In this case, we have $b=\left(\mu,\Sigma\right)$, where $\mu\in\mathbb{R}^{p}$ and $\Sigma\in\mathbb{R}^{p\times p}$ (p.s.d.). By assuming this parametric form for $g$, the constraints for a probability density are inherently satisfied. At every time step, the density function $g_{b}(\beta)$ is completely determined by $b$. By using the quadratic loss, the optimization problem, according to Equation 2.1, is as follows:
@@
(2.4) $\operatorname*{arg\,min}_{\mu,\Sigma}\left\{~\mathbb{E}~(y-\Delta(x)\beta)^{2}+\lambda\mathbb{E}~\beta^{\top}\beta+\Psi(\Sigma)~\right\}$
@@
We now compute each term in the objective function 2.4. Let $z:=\Delta(x)$ and since $\beta_{p}\sim\mathcal{N}(\mu,\Sigma)$,
$(y-z^{\top}\beta)^{2}=y^{2}-2yz^{\top}\beta+(z^{\top}\beta)^{2},$
@@
$\mathbb{E}(y-z^{\top}\beta)^{2}$ $=$ $\mathbb{E}y^{2}-2yz^{\top}\mathbb{E}\beta+\mathbb{E}(z^{\top}\beta)^{2}.$
@@
$\mathbb{E}[\beta]=\mu,$
$\text{Var}(\beta)=\Sigma=\mathbb{E}[\beta\beta^{\top}]-\mathbb{E}[\beta]\mathbb{E}[\beta]^{\top}\implies\mathbb{E}[\beta\beta^{\top}]=\Sigma+\mu\mu^{\top}.$
@@
$\mathbb{E}~(z^{\top}\beta)^{2}$ $=$ $\mathbb{E}~z^{\top}\beta\beta^{\top}z$
$=$ $z^{\top}\mathbb{E}[\beta\beta^{\top}]z$
$=$ $z^{\top}(\Sigma+\mu\mu^{\top})z$
$=$ $z^{\top}\Sigma z+z^{\top}\mu\mu^{\top}z$
$=$ $z^{\top}\Sigma z+(z^{\top}\mu)^{2}.$
@@
$\mathbb{E}\left[(y-z^{\top}\beta)^{2}\right]=y^{2}-2yz^{\top}\mu+z^{\top}\Sigma z+(z^{\top}\mu)^{2}.$
@@
$\lambda\mathbb{E}\beta^{\top}\beta=\lambda\left(\text{tr}\Sigma+\mu^{\top}\mu\right).$
And the constraint term enforces that the estimated covariance matrix of $\beta$ is p.s.d. which can be represented using an indicator function:
@@
$\underbrace{y^{2}-2yz^{\top}\mu+z^{\top}\Sigma\mathbf{x}+(z^{\top}\mu)^{2}+\lambda\left(\text{tr}\Sigma+\mu^{\top}\mu\right)}_{f_{1}}+\underbrace{1\!\!1_{p.s.d}(\Sigma)}_{f_{2}}$
@@
$\nabla_{\mu}f_{1}(\mu,\Sigma)=2z(z^{\top}\mu-y)+2\lambda\mu,$
@@
$\nabla_{\Sigma}f_{1}(\mu,\Sigma)=zz^{\top}+\lambda I.$
The indicator function in $f_{2}$ is non-smooth and the proximal operator is reduced to a projection onto a p.s.d. cone, following [1],
@@
Therefore, at the end of every time step, we update the parameters of $g_{t+1}$ based on the observed data $(z_{t},y_{t})$:
(2.5) $\mu^{t+1}=\mu^{t}-\eta_{t}\nabla_{\mu^{t}}f_{1}(\mu^{t},\Sigma^{t})=\mu^{t}-\eta_{t}\left(2z_{t}\left(z_{t}^{\top}\mu^{t}-y_{t}\right)+2\lambda\mu^{t}\right),$
$\widetilde{\Sigma}^{t+1}=\Sigma^{t}-\eta_{t}\nabla_{\Sigma^{t}}f_{1}(\mu^{t},\Sigma^{t})=\Sigma^{t}-\eta_{t}\left(z_{t}z^{\top}_{t}+\lambda I\right),$
(2.6) $\Sigma^{t+1}=\text{proj}_{K}\left(\widetilde{\Sigma}^{t+1}\right)$
@@
The prediction for the next output at step $t+1$, given the observed feature vector $z_{t+1}$, is computed as
$\hat{y}_{t+1}=z_{t+1}^{\top}\hat{\beta}_{t},\text{ where }\hat{\beta}_{t}=\mu^{t}$
and the corresponding $100(1-\alpha)\%$ prediction probability interval is given by
$\hat{y}_{t+1}\pm z_{\alpha/2}\sqrt{\widehat{Var}(\beta)_{t}},\text{ where }\widehat{Var}(\beta)_{t}=z_{t+1}^{\top}\Sigma^{t}z_{t+1}\text{ and }0<\alpha<1.$
2.2. Randomized SINDy for binary outcome
In a binary classification setting where $Y\in\{0,1\}$ and $z:=\Delta(x)\in\mathbb{R}^{p}$, we can adapt the previous example by utilizing the logistic function such that $h(z):=\sigma(z^{\top}\beta)$ represents the probability of predicting the output label $Y=1$:
$P(Y=1\mid z,\beta)=\sigma(z^{\top}\beta)=\dfrac{1}{1+\exp\{-z^{\top}\beta\}}.$